Solve: 10x^4 - 17x^2 + 3 = 0
Answer by mananth(16946)

10x^4 - 17x^2 + 3 = 0
10x^4-15x^2-2x^2+3=0
5x^2(2x^2-3)-1(2x^2-3)=0
(2x^2-3)((5x^2-1)=0
..
2x^2=3
x^2=3/2
x= + / - sqrt 3/2
..
OR
5x^2-1=0
5x^2=1
x^2=1/5
x=+ /- sqrt1/5